Grade 12th passAlgebra

Using laws of exponents, simplify and write the answer in exponential form:
( 220 ÷ 215 ) x 23

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

By the rule of dividing the powers with same base = am÷ an= am – n
( 220 ÷ 215)can be simplified as,
= (2)20 – 15
= 25
Then,
By the rule of multiplying the powers with same base = am × an = am + n
25 × 23 can be simplified as,
= (2)5+3
= 28

Thanks